OpenStudy (anonymous):

the smallest angle in a trangle 1/9 as large as the largest angle the third angle is eight times the smallest angle

OpenStudy (amistre64):

s+m+l = 180 :) s = l/9 m = 8s = 8(l/9) l + 8l/9 + l/9 = 180

OpenStudy (amistre64):

l (1 + 8/9 + 1/9) = 180 l = 180/2 l = 90

OpenStudy (amistre64):

s = 90/9 = 10 m = 80 small = 10, medium = 80, and large = 90

OpenStudy (anonymous):

thank you u rock